The energy of an excited H-atom is -3.4eV.Calculate the angular momentum of electron

Grade:11

1 Answers

Aarti Gupta
8 years ago
Given--
Energy of excited H-atom = – 3.4eV
Angular momentum of e- = ?
Solution--
We know that energy of an e- in nth orbit of hydrogen is given as--
En = – 13.6/n2 eV/atom
--13.6/n2 = – 3.4
therefore,
n2 = 13.6/3.4 = 4
so, n = 2
Now angular momentum of e- = nh/2\Pi = 2h/2\Pi
angular momentum = h/\Pi
Thus angular momentum of e- is h/pi